OpenStudy (anonymous):

find the equation of the normal to y=(2x-1)(3x+5)at the point (1,8). give your answer in the form ax+by+c=0, where a,b and c are integers

OpenStudy (phi):

did you find the derivative yet? once you do, set it equal to 13 (the gradient or slope) solve for x (2 values) then use x in the original to find the y so you get (x,y) for the answer

OpenStudy (anonymous):

yh i got 12x+7

OpenStudy (phi):

now? there are 3 terms with x in y=2x^3-5x^2+9x-1 term by term what do you get for the derivative?

OpenStudy (anonymous):

wait sorry i typed up the wrong question

OpenStudy (anonymous):

i've edited it now

OpenStudy (phi):

yes, I get dy/dx = 12x+7 (two ways to do it: use the product rule, or multiply it out then do term by term) Then want equation of the normal to y=(2x-1)(3x+5)at the point (1,8). normal means perpendicular. The slope of the tangent is dy/dx evaluated at x=1 the slope of the normal is -1/m first, find dy/dx when x=1

OpenStudy (anonymous):

19

OpenStudy (anonymous):

-1/19

OpenStudy (phi):

now what is the slope of the normal line? -1/19 last step y= mx + b with m= -1/19 x=1 and y=8 and solve for b

OpenStudy (anonymous):

152/19?

OpenStudy (phi):

I get 153/19

OpenStudy (anonymous):

oh right yh sorry i forgot to add on 1

OpenStudy (phi):

8+ 1/19 = (152+1)/19 = 153/19 not quite done, they want ax+by+c=0, where a,b and c are integers we have y= (-1/19)x + 153/19 mult by 19 19y = -x +153 move the stuff from the right to the left side

OpenStudy (anonymous):

x+19y-153=0?